### v1

# Existence of attracting periodic orbits in 3-dimensional strongly 2-cooperative systems

Abstract
The flow of a $k$-cooperative system maps the set of vectors with up to $(k-1)$ sign variations to itself. In particular, $1$-cooperative systems are just cooperative systems. Strongly $2$-cooperative systems satisfy a strong Poincaré-Bendixson property: any bounded solution that evolves in a compact set containing no equilibria converges to a periodic orbit. For $3$-dimensional strongly $2$-cooperative nonlinear systems, we provide a sufficient condition that guarantees the existence of an invariant compact set in the state space that includes an attracting periodic orbit. We show that our theoretical results unify and generalize known results on the existence of a periodic solution in two well-known models in biochemistry, a 3D Goodwin oscillator model and the 3D Field-Noyes ordinary-differential-equation (ODE) model for the Belousov-Zhabotinskii reaction, while simplifying the proofs.
Index Terms:
Asymptotic analysis, compound matrices, sign variations, competitive systems, systems biology.
I Introduction
In cooperative systems, an increase in one of the state variables can never decrease the derivative of another state variable: the state variables “cooperate” with one another. The flow of such systems maps the non-negative orthant to itself. More precisely, the flow maps the set of vectors with zero sign variations to itself. Cooperative systems have many special asymptotic properties: Hirsch’s quasi-convergence theorem asserts that in a strongly cooperative system almost every bounded solution converges to an equilibrium point; also, cooperative systems cannot have an attracting periodic orbit [30]. Cooperative systems (as well as interconnections of cooperative systems) and their generalization to cooperative control systems [2] have found numerous applications in fields including systems biology and chemistry [3, 7, 8, 18, 21, 31], dynamic neural networks [29], and social dynamics [28].
A recent generalization of nonlinear cooperative systems is called $k$-cooperative systems [35]. A system is $k$-cooperative if its flow maps the set of vectors with up to $k-1$ sign variations to itself. Just like cooperativity, $k$-cooperativity depends on the sign structure of the system Jacobian, and can be inferred even when the exact values of certain parameters are unknown. $1$-cooperative systems are cooperative systems, whereas $(n-1)$-cooperative systems (where $n$ is the dimension of the system) are, up to a coordinate transformation, competitive systems [35]. Another important case is strongly $2$-cooperative systems, as these satisfy a strong Poincaré-Bendixson property. This implies that if the closure of a trajectory of a bounded solution contains no equilibrium points, then that solution converges to a periodic orbit [35].
It is interesting to identify cases where a bounded solution does not converge to an equilibrium, but to a periodic orbit, implying that the system admits at least one attracting periodic solution, thereby allowing for oscillatory behavior. For example, in the context of systems biology Goodwin [14] states that: “It is of fundamental importance to an understanding of cellular organization whether or not the dynamic activity of molecular control processes involves oscillatory behavior.” The analysis cannot be based on the theory of cooperative systems, because the existence of an attracting periodic orbit automatically implies that the system is not cooperative.
Here, we consider 3-dimensional 2-cooperative systems. Our main result is a new simple sufficient condition for the existence of an attracting periodic orbit. The analysis is based on combining: (1) the strong Poincaré-Bendixson property of strongly $2$-cooperative systems; (2) the spectral properties of Jacobians of $2$-cooperative systems; and (3) an idea of Rauch [26] for constructing an invariant set $\Omega$ for a specific 3D model of a non-linear electric circuit such that $\Omega$ does not include an equilibrium.
We show that our theoretical results unify and generalize results on two specific non-linear 3D systems: a 3D Goodwin model and Field–Noyes 3D ODE model for the famous Belousov–Zhabotinskii reaction. The existence of a periodic solution is proven for the former model by Tyson [33] and for the latter model by Hastings and Murray [16]. However, in both cases, no information is provided on the stability of the periodic solution. Moreover, these papers rely on an application of the Schauder fixed-point theorem for proving existence of a periodic orbit, which is preceded by a fairly complicated preliminary analysis.
Our main result employs much simpler techniques than both [33] and [16] for proving existence of a periodic orbit and further implies that each of these systems admits at least one attracting periodic orbit. This is a fundamental extension of [33, 16] because, from an application viewpoint, a periodic orbit $\gamma$ which is not attracting is less interesting, as it yields no guarantees on oscillatory behavior, meaning that only solutions emanating from an initial condition precisely on the periodic orbit, i.e. $x(0)\in\gamma$, are guaranteed to induce oscillations.
We mention in passing that reassessing these papers from the point of view of $2$-cooperative systems shows that the proof in [16] is, in fact, incomplete. [^1] In particular, analyzing the sign structure of the sets $B_{i}$ in Eq. (20) in [16] and using the sign variation properties of 2-cooperative systems shows that the property described in this equation is wrong.
The remainder of this note is organized as follows. The next section reviews definitions and results that will be used subsequently. Section III details our main result. Section IV shows that our main result implies the existence of an attracting periodic orbit in a 3D Goodwin model and in the Field–Noyes 3D ODE model for the Belousov–Zhabotinskii reaction. The final section concludes, and describes several possible directions for further research.
II Notation and Preliminary Results
We denote vectors and matrices by lowercase and uppercase letters, respectively. For two vectors $x,y\in{\mathbb{R}}^{n}$, we write $x\geq y$ if $x_{i}\geq y_{i}$ for all $i$. The non-negative orthant in ${\mathbb{R}}^{n}$ is
${\mathbb{R}}^{n}_{\geq 0}:=\{x\in{\mathbb{R}}^{n}\,|\,x\geq 0\}.$
The transpose and determinant of a matrix $A$ are denoted by $A^{\top}$ and $\det(A)$ respectively, while $I_{n}$ denotes the $n\times n$ identity matrix. A square matrix $A$ is called Hurwitz if all its eigenvalues have a negative real part, unstable if it admits an eigenvalue with a positive real part, and Metzler if all its off-diagonal entries are non-negative. A matrix $\bar{A}$ is a sign matrix if every entry of $\bar{A}$ is either * (standing for “don’t care”), $\leq 0$, or $\geq 0$. A time-varying matrix $A(t)$ has the sign pattern $\bar{A}$ if the following three properties hold at all times $t$:
  1. 1.
    $a_{ij}(t)=0$ for all indices $i,j$ such that $\bar{a}_{ij}=0$;
  2. 2.
    $a_{ij}(t)\geq 0$ for all indices $i,j$ such that $\bar{a}_{ij}$ equals $\geq 0$;
  3. 3.
    $a_{ij}(t)\leq 0$ for all indices $i,j$ such that $\bar{a}_{ij}$ equals $\leq 0$,
while there is no restriction on $a_{ij}(t)$ when $\bar{a}_{ij}=*$. Given a set $S$, $\operatorname{int}(S)$ denotes its interior, and by $|S|$ its cardinality.
As mentioned earlier, the flow of a cooperative system maps ${\mathbb{R}}^{n}_{\geq 0}$ to ${\mathbb{R}}^{n}_{\geq 0}$, and it also maps ${\mathbb{R}}^{n}_{\leq 0}:=-({\mathbb{R}}^{n}_{\geq 0})$ to ${\mathbb{R}}^{n}_{\leq 0}$. In other words, the flow maps the set of vectors with zero sign variations to itself. A $k$-cooperative system maps the set of vectors with up to $k-1$ sign variations to itself.
Sign variations in a vector. Let $\sigma(x)$ denote the number of sign variations in a vector $x\in{\mathbb{R}}^{n}$ with no zero entries; for example, $\sigma\left(\begin{bmatrix}6.3&-\pi&1\end{bmatrix}^{\top}\right)=2$. The theory of totally positive matrices (see, e.g,. [9, 12, 25]) offers two useful generalizations of $\sigma(\cdot)$ to vectors that may include zero entries.
Definition 1.
Given $x\in{\mathbb{R}}^{n}\setminus\{0\}$, let $y$ denote the vector $x$ after deleting all its zero entries. Define
$s^{-}(x):=\sigma(y),$
and
$s^{+}(x):=\max_{z\in\mathcal{S}_{x}}\sigma(z),$
where $\mathcal{S}_{x}$ is the set of vectors obtained from $x$ by replacing each zero entry by either $+1$ or $-1$.
For example, for $x=\begin{bmatrix}-1&0&0&1&2\end{bmatrix}^{\top}$, we have $s^{-}(x)=\sigma(\begin{bmatrix}-1&1&2\end{bmatrix}^{\top})=1$, whereas $s^{+}(x)=\sigma(\begin{bmatrix}-1&1&-1&1&2\end{bmatrix}^{\top})=3$. If $x\in{\mathbb{R}}^{n}$ has no zero entries, then $s^{-}(x)=s^{+}(x)=\sigma(x)$. For the zero vector $0\in{\mathbb{R}}^{n}$, we define $s^{-}(0):=0$, and $s^{+}(0):=n-1$. Then
$0\leq s^{-}(x)\leq s^{+}(x)\leq n-1,\text{ for all }x\in{\mathbb{R}}^{n}.$
We will be particularly interested in vectors $x\in{\mathbb{R}}^{n}$ such that $s^{+}(x)\leq 1$. This always holds for $x\in{\mathbb{R}}^{2}$, so for the rest of this section we consider ${\mathbb{R}}^{n}$ with $n\geq 3$. Then $s^{+}(x)\leq 1$ implies in particular that $x_{1}^{2}+x_{n}^{2}\neq 0$.
Fix $x,y\in{\mathbb{R}}^{n}$, with $x\not=y$, so that $z:=x-y\not=0$; if $s^{+}(z)\leq 1$, then also the projection to the $2$-dimensional vector $\begin{bmatrix}z_{1}&z_{n}\end{bmatrix}^{\top}=\begin{bmatrix}x_{1}-y_{1}&x_{n}-y_{n}\end{bmatrix}^{\top}\not=0$.
The set of vectors with up to $k-1$ sign variations can be defined using either $s^{-}(x)$ or $s^{+}(x)$.
Definition 2.
Given $k\in\{1,\dots,n\}$, let
$P^{k}_{-}$ :=$:=\{x\in{\mathbb{R}}^{n}\ |\ s^{-}(x)\leq k-1\},$
$P^{k}_{+}$ :=$:=\{x\in{\mathbb{R}}^{n}\ |\ s^{+}(x)\leq k-1\}.$
For example, $P^{1}_{-}={\mathbb{R}}^{n}_{\geq 0}\cup{\mathbb{R}}^{n}_{\leq 0}$, and $P^{1}_{+}=\operatorname{int}({\mathbb{R}}^{n}_{\geq 0}\cup{\mathbb{R}}^{n}_{\leq 0})$. More generally, it can be shown [35] that $P^{k}_{-}$ is closed, and that $P^{k}_{+}=\operatorname{int}(P^{k}_{-})$ for all $k$. By definition, these sets have a nested structure
$P^{1}_{-}\subset P^{2}_{-}\subset\dots\subset P^{n}_{-}={\mathbb{R}}^{n},$
$P^{1}_{+}\subset P^{2}_{+}\subset\dots\subset P^{n}_{+}={\mathbb{R}}^{n}.$
A set $C\subseteq{\mathbb{R}}^{n}$ is a cone if $x\in C$ implies that $\alpha x\in C$ for all $\alpha\in{\mathbb{R}}_{\geq 0}$. It is straightforward to verify that $P^{k}_{-}$ is a cone for any $k$. However, it is not a convex cone. For example, the vectors $y^{1}:=\begin{bmatrix}-2&1&1\end{bmatrix}^{\top}\in P^{2}_{-}$ and $y^{2}:=\begin{bmatrix}1&1&-2\end{bmatrix}^{\top}\in P^{2}_{-}$, but $\frac{1}{2}y^{1}+\frac{1}{2}y^{2}=\frac{1}{2}\begin{bmatrix}-1&2&-1\end{bmatrix}^{\top}\not\in P^{2}_{-}$.
To understand the geometry of $P^{k}_{-}$, recall that a closed set $C\subseteq{\mathbb{R}}^{n}$ is called a cone of rank $k$ if (1) $x\in C$ implies that $\alpha x\in C$ for all $\alpha\in{\mathbb{R}}$; and (2) $C$ contains a linear space of dimension $k$, and no linear space of a higher dimension [17]. For example, $P^{1}_{-}={\mathbb{R}}^{n}_{+}\cup{\mathbb{R}}^{n}_{-}$ is a cone of rank $1$. A cone $C$ of rank $k$ is called $k$-solid if there exists a $k$-dimensional linear subspace $V$ such that $V\setminus\{0\}\subseteq\operatorname{int}(C)$. For all $k\in\{1,\dots,n\}$, the set $P^{k}_{-}$ is a cone of rank $k$ that is $k$-solid [35].
Roughly speaking, if a dynamical system admits a $k$-solid cone as an invariant set, and if its trajectories can be projected in a one-to-one way on the $k$-dimensional linear subspace $V$ then its trajectories “behave” as those of a $k$-dimensional system. In particular, if $k=2$, then the trajectories “behave” as those of a planar dynamical system [27], thereby allowing to apply analysis techniques that are suitable for planar systems.
$k$-positive linear dynamical systems. We now recall (strong) $k$-positivity for linear time-varying systems.
Definition 3.
The linear time-varying (LTV) system
$\dot{x}(t)=A(t)x(t)$ (1)
is $k$-positive if its flow maps $P^{k}_{-}$ to itself, and strongly $k$-positive if its flow maps $P^{k}_{-}\setminus\{0\}$ to $P^{k}_{+}$.
Under the assumption that $t\to A(t)$ is continuous, the linear system (1) is $1$-positive (i.e., positive) iff $A(t)$ is Metzler for all $t$, that is, iff, for all $t$, $A(t)$ has the sign pattern
$\bar{A}_{1}:=\small\begin{bmatrix}*&\geq 0&\geq 0&\dots&\geq 0&\geq 0\\ \geq 0&*&\geq 0&\dots&\geq 0&\geq 0\\ \geq 0&\geq 0&*&\dots&\geq 0&\geq 0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ \geq 0&\geq 0&\geq 0&\dots&*&\geq 0\\ \geq 0&\geq 0&\geq 0&\dots&\geq 0&*\end{bmatrix}.$
The linear system (1) is $2$-positive iff, for all $t$$A(t)$ has the sign pattern
$\bar{A}_{2}:=\small\begin{bmatrix}*&\geq 0&0&\dots&0&\leq 0\\ \geq 0&*&\geq 0&\dots&0&0\\ 0&\geq 0&*&\dots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\dots&*&\geq 0\\ \leq 0&0&0&\dots&\geq 0&*\end{bmatrix},$ (2)
while it is strongly $2$-positive if, in addition, the matrix $A(t)$ is irreducible for almost all $t$ [35]. Since the upper-right and lower-left corner entries of $A(t)$ may take negative values, $A(t)$ is not necessarily Metzler.
Remark 1.
The system $\dot{x}(t)=\bar{A}_{2}x(t)$ may be interpreted as a multiagent system with agents $x_{1},\dots,x_{n}$ interconnected by bidirectional links with a ring topology. Each $x_{i}$, $i\in\{2,\dots,n-1\}$ receives arbitrary signed feedback from itself, and non-negative feedback from its left- and right-neighbours $x_{i-1}$ and $x_{i+1}$. Agent $x_{1}$ receives arbitrary signed feedback from itself, non-negative feedback from its right neighbour $x_{2}$ and non-positive feedback from its “cyclic left” neighbour $x_{n}$. Agent $x_{n}$ receives arbitrary signed feedback from itself, non-negative feedback from its left neighbour $x_{n-1}$ and non-positive feedback from its “cyclic right” neighbour $x_{1}$.
$k$-cooperative nonlinear dynamical systems. A nonlinear dynamical system is called $k$-cooperative if its associated variational system, which is a linear time-varying system, is $k$-positive for all times $t\geq 0$ and all points $x$ in the state space. Consider the time-invariant non-linear system
$\dot{x}=f(x),$ (3)
whose solutions evolve on a convex state space $\Omega\subseteq{\mathbb{R}}^{n}$. Assume that $f$ is $C^{1}$, with Jacobian $J(x):=\frac{\partial}{\partial x}f(x)$, and that for all initial conditions $a\in\Omega$ the system admits a unique solution $x(t,a)\in\Omega$ for all $t\geq 0$. For two initial conditions $a,b\in\Omega$, let $z(t):=x(t,a)-x(t,b)$. Then
$\dot{z}$ $=M(t)z,$ (4)
where
$M(t)$ $:=\int_{0}^{1}J\big{(}rx(t,a)+(1-r)x(t,b)\big{)}\mathop{}\!\mathrm{d}r.$
Note that (4) is an LTV, and that, if $J(x)$ satisfies some sign pattern condition for all $x\in\Omega$, then $M(t)$ has the same sign pattern for all $t\geq 0$, since sign pattern conditions are preserved under integration.
Definition 4.
The nonlinear system (3) is (strongly) $k$-cooperative if the associated variational system (4) is (strongly) $k$-positive for all $a,b\in\Omega$ and all $t\geq 0$.
For example, the system is $1$-cooperative if the variational system (4) is $1$-positive, that is, if $J(x)$ is Metzler for all $x$. Thus $1$-cooperativity is just cooperativity.
Remark 2.
Suppose that 00 is an equilibrium of (3). Fix $a\in\Omega$, with $a\not=0$, and define $z(t):=x(t,a)-x(t,0)=x(t,a)$. If the non-linear system (3) is strongly $k$-cooperative, it follows from Definitions 3 and 4 and from Eq. (4) that, if $s^{-}(x(t,a))\leq k-1$ for some time $t\geq 0$, then $s^{+}(x(\tau,a))\leq k-1$ for all $\tau>t$.
For an initial condition $a\in\Omega$, let $\omega(a)$ denote the omega limit set of the solution of (3) emanating from $x(0)=a$, namely, the set of all points $y\in\Omega$ for which there exists a sequence $(t_{n})_{n\in\mathbb{N}}$ in ${\mathbb{R}}$, with $\lim_{n\to\infty}t_{n}=+\infty$, such that $\lim_{n\to\infty}x(t_{n},a)=y$; see e.g. [32, p. 193].
Definition 5.
The dynamical system (3), having the set of equilibria $\mathcal{E}$, satisfies the strong Poincaré-Bendixson property if, for any bounded solution $x(t,a)$, with $a\in\Omega$, it holds that
$\omega(a)\cap\mathcal{E}=\emptyset\implies\omega(a)\ \text{is a periodic orbit}.$
Intuitively, this suggests that all solutions behave like the solutions of a planar dynamical system.
Since a strongly $2$-cooperative system satisfies the strong Poincaré-Bendixson property [35], establishing strong $2$-cooperativity provides important information on the asymptotic behaviour of the non-linear system.
The analysis of sign patterns guaranteeing $k$-positivity and $k$-cooperativity relies on the theory of compound matrices (see, e.g., [4]), a fundamental tool also for $k$-contractive [37], $\alpha$-contractive [38], and totally positive differential systems [20]. A sign pattern related to (2) appears in the seminal work by Mallet-Paret and Smith on monotone cyclic feedback systems [19] (see also [5, 10, 34] and the references therein for more recent results).
Instability of a $3\times 3$ strongly $2$-cooperative matrix. A matrix $A\in{\mathbb{R}}^{n\times n}$ with a strongly $k$-positive sign pattern has special spectral properties [1]. Here, we state these properties in the special case with $n=3$, $k=2$, and $A$ is unstable, that is, it admits an eigenvalue with a positive real part.
Lemma 1.
Given $A\in{\mathbb{R}}^{3\times 3}$, consider its characteristic polynomial
$P_{A}(s):=\det(sI_{3}-A)=s^{3}+a_{2}s^{2}+a_{1}s+a_{0}.$
If the free coefficient $a_{0}$ is positive and $A$ is unstable, then $A$ admits one negative real eigenvalue, and two eigenvalues with a positive real part. If, in addition, the system $\dot{x}=Ax$ is strongly $2$-positive, then the eigenvector $\zeta\in{\mathbb{R}}^{3}$ corresponding to the negative real eigenvalue has sign pattern either $\begin{bmatrix}+&-&+\end{bmatrix}^{\top}$ or $\begin{bmatrix}-&+&-\end{bmatrix}^{\top}$, so
$s^{-}(\zeta)=2.$ (5)
Consider the system $\dot{x}=Ax$ and let $x(t)$ be a solution. Lemma 1 implies that $x(t)$ converges to the origin iff $x(t)\in\operatorname{span}(\zeta)$ for any $t$, meaning that $\zeta$ is the only “direction of convergence” to the equilibrium for the system.
Proof:
Let $\lambda_{i}$, $i=1,2,3$, denote the eigenvalues of $A$. Since $P_{A}(s)$ has real coefficients and is of odd degree, $A$ has at least one real eigenvalue. Moreover, since $A$ is unstable, one of its eigenvalues has a positive real part. Since $\lambda_{1}\lambda_{2}\lambda_{3}=-a_{0}<0$, matrix $A$ must admit one negative real eigenvalue, and the other two eigenvalues must be either both positive real, or a complex-conjugate pair with a positive real part.
The determinant of a $2\times 2$ submatrix of a matrix $B\in{\mathbb{R}}^{n\times n}$ is a $2\times 2$ minor of $B$. If $\dot{x}=Ax$ is strongly 2-positive, then all the $2\times 2$ minors of $\exp(A)$ are positive [35]. The eigenvalues of $\exp(A)$ are $\exp(\lambda_{i})$, $i=1,2,3$. Order the $\lambda_{i}$s so that $|\exp(\lambda_{1})|\geq|\exp(\lambda_{2})|\geq|\exp(\lambda_{3})|$. Then, [1, Thm. 2] implies that the product $\exp(\lambda_{1})\exp(\lambda_{2})$ is real and positive, and that the eigenvalues of $\exp(A)$ satisfy the spectral gap condition
$|\exp(\lambda_{2})|>|\exp(\lambda_{3})|.$ (6)
Hence, $\lambda_{3}$ is the real and negative eigenvalue. Using [1, Thm. 2] again gives that the eigenvector $\zeta$ of matrix $\exp(A)$ corresponding to the eigenvalue $\exp(\lambda_{3})$ satisfies (5), which completes the proof. ∎
The next simple example demonstrates Lemma 1.
Example 1.
Let $A:=\begin{bmatrix}-1&0&-9\\ 1&-1&0\\ 0&1&-1\end{bmatrix}$. The system $\dot{x}=Ax$ is strongly $2$-positive, because $A$ is irreducible and has the sign pattern $\bar{A}_{2}$. We can compute numerically $\exp(A)=\begin{bmatrix}-0.1433&-1.4137&-2.1219\\ 0.2358&-0.1433&-1.4137\\ 0.1571&0.2358&-0.1433\end{bmatrix}$ (all numerical values in this paper are to four-digit accuracy), as well as its $2\times 2$ minors: ($0.3538$, $0.7028$, $1.6946$, $0.1883$, $0.3538$, $0.7028$, $0.0781$, $0.1883$, $0.3538$), which are all positive.
The characteristic polynomial of matrix $A$ is $P_{A}(s)=\det(sI_{3}-A)=s^{3}+3s^{2}+3s+10$, with a positive free coefficient. Its eigenvalues are $\lambda_{1}=\mu+\jmath\eta$,  $\lambda_{2}=\mu-\jmath\eta$, and $\lambda_{3}=-3.0801$, where $\mu=0.0400$, $\eta=1.8014$ and $\jmath$ is the imaginary unit. Hence, $A$ is unstable. The eigenvector corresponding to the negative real eigenvalue $\lambda_{3}$ is $\zeta=\begin{bmatrix}-0.8823&0.4242&-0.2039\end{bmatrix}^{\top}$, with $s^{-}(\zeta)=2$.
III Existence of an Attracting Periodic Orbit
We can now state and prove our main result on the existence of an attracting periodic orbit in strongly $2$-cooperative 3D systems.
Theorem 1.
Consider the non-linear time-invariant system
$\dot{x}=f(x),\quad x\in{\mathbb{R}}^{3}.$ (7)
Suppose that $f\in C^{1}$, and let $J(x):=\frac{\partial}{\partial x}f(x)$ denote the Jacobian of the vector field. Suppose that the system is strongly 2-cooperative, and that its trajectories evolve in the closed box
=${\mathcal{B}}=\{x\in{\mathbb{R}}^{3}\ |\ \underline{x}\leq x\leq\overline{x}\},$
for some $\underline{x},\overline{x}\in{\mathbb{R}}^{3}$ with $\underline{x}\leq\overline{x}$. Suppose also that ${\mathcal{B}}$ contains a unique equilibrium $e\in\operatorname{int}({\mathcal{B}})$ that is unstable, and that the free coefficient in the characteristic polynomial $P(s)=\det(sI_{3}-J(e))$ is positive. Partition ${\mathcal{B}}$ into eight closed sub-boxes:
${\mathcal{B}}_{1}:=\{x\in{\mathcal{B}}\,|\,x_{1}\leq e_{1},\;x_{2}\leq e_{2},\;x_{3}\leq e_{3}\},$
${\mathcal{B}}_{2}:=\{x\in{\mathcal{B}}\,|\,x_{1}\geq e_{1},\;x_{2}\leq e_{2},\;x_{3}\leq e_{3}\},$
${\mathcal{B}}_{3}:=\{x\in{\mathcal{B}}\,|\,x_{1}\geq e_{1},\;x_{2}\geq e_{2},\;x_{3}\leq e_{3}\},$
${\mathcal{B}}_{4}:=\{x\in{\mathcal{B}}\,|\,x_{1}\geq e_{1},\;x_{2}\geq e_{2},\;x_{3}\geq e_{3}\},$
${\mathcal{B}}_{5}:=\{x\in{\mathcal{B}}\,|\,x_{1}\leq e_{1},\;x_{2}\geq e_{2},\;x_{3}\geq e_{3}\},$
${\mathcal{B}}_{6}:=\{x\in{\mathcal{B}}\,|\,x_{1}\leq e_{1},\;x_{2}\leq e_{2},\;x_{3}\geq e_{3}\},$
${\mathcal{B}}_{7}:=\{x\in{\mathcal{B}}\,|\,x_{1}\geq e_{1},\;x_{2}\leq e_{2},\;x_{3}\geq e_{3}\},$
${\mathcal{B}}_{8}:=\{x\in{\mathcal{B}}\,|\,x_{1}\leq e_{1},\;x_{2}\geq e_{2},\;x_{3}\leq e_{3}\}.$
Then ${\mathcal{B}}_{16}:={\mathcal{B}}_{1}\cup\dots\cup{\mathcal{B}}_{6}$ is an invariant set, and for any initial condition $x(0)\in({\mathcal{B}}_{16}\setminus\{e\})$, the solution of (7) converges to a (non-trivial) periodic orbit.
This result implies of course that ${\mathcal{B}}_{16}$ includes at least one attracting periodic orbit. Since cooperative (that is, 1-cooperative) systems do not admit attracting periodic orbits, it is clear that the result of Thm. 1 cannot be derived using the theory of cooperative systems.
Proof:
Define $z_{i}(t):=x_{i}(t)-e_{i}$, $i=1,2,3$. Then,
$\dot{z}=f(z+e),$ (8)
and the trajectories of this system evolve in the shifted closed box ~$\tilde{\mathcal{B}}:={{\mathcal{B}}}-e$. In the $z$-coordinates, the unique equilibrium is at the origin, and the sub-boxes defined in the theorem statement become
~$\tilde{\mathcal{B}}_{1}=\{z\in\tilde{\mathcal{B}}\ |\ z_{1}\leq 0,\;z_{2}\leq 0,\;z_{3}\leq 0\},$
~$\tilde{\mathcal{B}}_{2}=\{z\in\tilde{\mathcal{B}}\ |\ z_{1}\geq 0,\;z_{2}\leq 0,\;z_{3}\leq 0\},$
~$\tilde{\mathcal{B}}_{3}=\{z\in\tilde{\mathcal{B}}\ |\ z_{1}\geq 0,\;z_{2}\geq 0,\;z_{3}\leq 0\},$
~$\tilde{\mathcal{B}}_{4}=\{z\in\tilde{\mathcal{B}}\ |\ z_{1}\geq 0,\;z_{2}\geq 0,\;z_{3}\geq 0\},$
~$\tilde{\mathcal{B}}_{5}=\{z\in\tilde{\mathcal{B}}\ |\ z_{1}\leq 0,\;z_{2}\geq 0,\;z_{3}\geq 0\},$
~$\tilde{\mathcal{B}}_{6}=\{z\in\tilde{\mathcal{B}}\ |\ z_{1}\leq 0,\;z_{2}\leq 0,\;z_{3}\geq 0\},$
~$\tilde{\mathcal{B}}_{7}=\{z\in\tilde{\mathcal{B}}\ |\ z_{1}\geq 0,\;z_{2}\leq 0,\;z_{3}\geq 0\},$
~$\tilde{\mathcal{B}}_{8}=\{z\in\tilde{\mathcal{B}}\ |\ z_{1}\leq 0,\;z_{2}\geq 0,\;z_{3}\leq 0\}.$
For subsequent arguments, it is useful to list the number of sign variations of vectors in these sub-boxes:
$z\in\tilde{\mathcal{B}}_{1}$implies $s^{-}(z)=0$,
$z\in\tilde{\mathcal{B}}_{2}$implies $s^{-}(z)\leq 1$,
$z\in\tilde{\mathcal{B}}_{3}$implies $s^{-}(z)\leq 1$,
$z\in\tilde{\mathcal{B}}_{4}$implies $s^{-}(z)=0$,
$z\in\tilde{\mathcal{B}}_{5}$implies $s^{-}(z)\leq 1$,
$z\in\tilde{\mathcal{B}}_{6}$implies $s^{-}(z)\leq 1$,
$z\in\tilde{\mathcal{B}}_{7}$implies $s^{+}(z)=2$,
$z\in\tilde{\mathcal{B}}_{8}$implies $s^{+}(z)=2$.
Since the system is strongly 2-cooperative, Remark 2 implies that
$z(t)\not=0\text{ and }s^{-}(z(t))\leq 1\text{ for some }t\geq 0$
$\implies s^{+}(z(\tau))\leq 1\text{ for all }\tau>t.$ (9)
Hence, ~$\tilde{\mathcal{B}}_{16}:=\tilde{\mathcal{B}}_{1}\cup\dots\cup\tilde{\mathcal{B}}_{6}$ is an invariant set for (8). Note that the equilibrium 00 is contained in ~$\tilde{\mathcal{B}}_{16}$. Hence, a priori, a solution evolving in ~$\tilde{{\mathcal{B}}}_{16}$ may still converge to the equilibrium. Our goal now is to build an invariant set for the dynamics (8) by cutting out from ~$\tilde{\mathcal{B}}_{16}$ a cylinder that contains the origin.
To do this, we first analyze the behaviour near the equilibrium. Let $A$ denote the Jacobian of (8) at the origin. Since the origin is unstable, $A$ is unstable, and Lemma 1 implies that $A$ has two eigenvalues $\lambda_{1},\lambda_{2}$ with a positive real part, and one real eigenvalue $\lambda_{3}<0$. Furthermore, the shifted eigenvector
$\tilde{\zeta}:=\zeta-e$
corresponding to $\lambda_{3}$ satisfies
$(\operatorname{span}(\tilde{\zeta})\setminus\{0\})\cap\tilde{{\mathcal{B}}}_{16}\subseteq\operatorname{int}(\tilde{\mathcal{B}}_{7}\cup\tilde{\mathcal{B}}_{8}).$
In particular, each nonzero solution $z(t)$ of (8) that converges to the origin must satisfy $\lim_{t\to\infty}\operatorname{dist}\left(z(t),\operatorname{span}(\tilde{\zeta})\right)=0$.
In view of Lemma 1, there exists a non-singular $T\in\mathbb{R}^{3\times 3}$ such that
$T^{-1}AT=\small\begin{bmatrix}\lambda_{3}&0&0\\ 0&u_{1}&-v\\ 0&v&u_{2}\end{bmatrix},$ (10)
with $\lambda_{3}<0$, $u_{1},u_{2}>0$, where either (1) $v=0$, so $u_{1},u_{2}$ are two positive real eigenvalues; or (2) $u_{1}=u_{2}$, so $u_{1}\pm jv$ are a complex conjugate pair of unstable eigenvalues.
Let $\angle(a,b)$ denote the angle between the vectors $a,b\in{\mathbb{R}}^{3}$. Since $\tilde{\zeta}\in\operatorname{int}\left(\tilde{{\mathcal{B}}}_{7}\cup\tilde{{\mathcal{B}}}_{8}\right)$, there exists $\xi\in(0,1)$ such that, for all $z\in{\mathcal{B}}\setminus\{0\}$, we have
$|\cos$ $(\angle(T^{-1}z,T^{-1}\tilde{\zeta}))|>1-\xi$
$\implies T^{-1}z\in\operatorname{int}(T^{-1}(\tilde{{\mathcal{B}}}_{7}\cup\tilde{{\mathcal{B}}}_{8})).$
By (10), $T^{-1}\tilde{\zeta}=\begin{bmatrix}1&0&0\end{bmatrix}^{\top}$, so letting $q:=T^{-1}z$ yields
$|\cos(\angle(q,\begin{bmatrix}1&0&0\end{bmatrix}^{\top}))|>1-\xi\implies q\in\operatorname{int}(T^{-1}(\tilde{{\mathcal{B}}}_{7}\cup\tilde{{\mathcal{B}}}_{8})),$ (11)
(see Fig. 1).
Figure 1: The red line is $\operatorname{span}(\tilde{\zeta})$, and the cubes are ~$\tilde{{\mathcal{B}}}_{7}$ and ~$\tilde{{\mathcal{B}}}_{8}$, intersecting at the origin. Left: Any $z\in\tilde{{\mathcal{B}}}$ such that $\angle(z,\tilde{\zeta})$ is sufficiently close to 00 or $\pi$ lies in the interior of the two cubes. Right: The invariant set is obtained by cutting out from ~$\tilde{{\mathcal{B}}}_{16}$ a cylinder around $\operatorname{span}(\tilde{\zeta})$. Thus, the invariant set has a positive distance from the equilibrium at the origin.
Define the state vector $q(t):=T^{-1}z(t)$. Then
$\dot{q}=h(q),$ (12)
with
$h(q)=\small\begin{bmatrix}\lambda_{3}&0&0\\ 0&u_{1}&-v\\ 0&v&u_{2}\end{bmatrix}q+\begin{bmatrix}g_{1}(q)\\ g_{2}(q)\\ g_{3}(q)\end{bmatrix},$ (13)
where the functions $g_{i}(q)$, $i=1,2,3$, are smooth on the compact set $T^{-1}\tilde{\mathcal{B}}$, and there exists $M>0$ such that
$\max_{q\in T^{-1}\tilde{\mathcal{B}}}\max_{i}\frac{\left|g_{i}(q)\right|}{|q|^{2}}\leq M.$ (14)
Define $V:{\mathbb{R}}^{2}\to{\mathbb{R}}_{\geq 0}$ by
$V(p_{1},p_{2}):=(u_{1}^{-1}p_{1}^{2}+u_{2}^{-1}p_{2}^{2})/2.$
We claim that there exists $\eta_{*}>0$ such that for all $0<\eta<\eta_{*}$ the closed set
~$\tilde{\mathcal{H}}_{\eta}:=\{q\in T^{-1}\tilde{\mathcal{B}}_{16}\,|\,V(q_{2},q_{3})\geq\eta\},$ (15)
is an invariant set of (12). To prove this, Let $\eta>0$ and fix an initial condition $q(0)\in\tilde{\mathcal{H}}_{\eta}$. Note that (12) implies that
$\dot{V}(q_{2}(t),q_{3}(t))$ $=u_{1}^{-1}q_{2}(t)\dot{q}_{2}(t)+u_{2}^{-1}q_{3}(t)\dot{q}_{3}(t)$
$=q_{2}^{2}(t)+q_{3}^{2}(t)+s(t),$ (16)
where $s(t):=(u_{2}^{-1}-u_{1}^{-1})vq_{2}(t)q_{3}(t)+u_{1}^{-1}q_{2}(t)g_{2}(q(t))+u_{2}^{-1}q_{3}(t)g_{3}(q(t))$. Since either $v=0$ (in the case of two real positive eigenvalues), or $u_{1}=u_{2}$ (in the case of a complex conjugate pair of unstable eigenvalues), we have
$s(t)=u_{1}^{-1}q_{2}(t)g_{2}(q(t))+u_{2}^{-1}q_{3}(t)g_{3}(q(t)).$
Since $q(t)\in T^{-1}\tilde{{\mathcal{B}}}_{16}$, Eq. (11) implies that
$|\cos(\angle(q(t),\begin{bmatrix}1&0&0\end{bmatrix}^{\top}))|\leq 1-\xi,$ (17)
that is, $\frac{q_{1}^{2}(t)}{|q(t)|^{2}}\leq(1-\xi)^{2}$, and since $\xi\in(0,1)$, this gives $|q(t)|^{2}\leq\frac{1}{1-(1-\xi)^{2}}(q_{2}^{2}(t)+q_{3}^{2}(t))$. Combining this with (14) yields
$|q_{i}(t)g_{i}(q(t))|$ $\leq|q_{i}(t)|\frac{|g_{i}(q(t))|}{|q(t)|^{2}}|q(t)|^{2}$
$\leq|q_{i}(t)|\frac{M}{1-(1-\xi)^{2}}(q_{2}^{2}(t)+q_{3}^{2}(t)),$ (18)
for $i\in\{2,3\}$. Thus,
$|s(t)|$ $\leq\frac{M}{1-(1-\xi)^{2}}\big{(}u_{1}^{-1}(q_{2}^{2}(t)|q_{2}(t)|+q_{3}^{2}(t)|q_{2}(t)|)$
$+u_{2}^{-1}(q_{2}^{2}(t)|q_{3}(t)|+q_{3}^{2}(t)|q_{3}(t)|)\big{)}.$
Combining this with (III) implies that there exists a sufficiently small $\eta_{*}>0$ such that, for all $0<\eta<\eta_{*}$ and $q(t)\in\tilde{\mathcal{H}}_{\eta}$, we have that $\dot{V}(q_{2}(t),q_{3}(t))\geq 0$, and thus ~$\tilde{\mathcal{H}}_{\eta}$ is indeed an invariant set of the dynamics. By the definition of ~$\tilde{\mathcal{H}}_{\eta}$, the equilibrium point 00 is not in ~$\tilde{\mathcal{H}}_{\eta}$ (see Fig. 1).
Summarizing, for some $\eta_{*}>0$ and any $0<\eta<\eta_{*}$, ~$\tilde{\mathcal{H}}_{\eta}$ is a closed invariant set, and for all $q(0)\in\tilde{\mathcal{H}}_{\eta}$ the solution $q(t)$ is bounded and keeps a positive distance from the equilibrium 00 (and thus from any equilibrium). Combining this with the strong Poincaré-Bendixson property of the system implies that any trajectory emanating from ~$\tilde{\mathcal{H}}_{\eta}$ converges to a periodic orbit. Since any $q(0)\in\tilde{{\mathcal{B}}}_{16}\setminus\{0\}$ belongs to $\cup_{0<\eta<\eta_{*}}\tilde{\mathcal{H}}_{\eta}$, this completes the proof of Thm. 1. ∎
IV Case Studies
We describe two case studies that demonstrate the application of Thm. 1 to well-known models in systems biology and chemistry.
IV-A 3-dimensional Goodwin oscillator model
The Goodwin oscillator is a classical biochemical control circuit that regulates enzyme or protein synthesis by negative end-product feedback [14]. Consider the 3D version of the model,
$\begin{cases}\dot{x}_{1}&=-\alpha x_{1}+\frac{1}{1+x_{3}^{m}},\\ \dot{x}_{2}&=-\beta x_{2}+x_{1},\\ \dot{x}_{3}&=-\gamma x_{3}+x_{2},\end{cases}$ (19)
where $\alpha,\beta,\gamma>0$, and $m$ is a positive integer.
As noted by Gonze and Ruoff [13]: “The three-variable Goodwin model (adapted by Griffith) can be seen as a core model for a large class of biological systems, ranging from ultradian to circadian clocks.”
The state space of (19) is $\Omega:={\mathbb{R}}^{3}_{\geq 0}$. Furthermore, Griffith [15] showed that all trajectories are bounded, as any trajectory emanating from $\Omega$ eventually enters into the closed box
${\mathcal{B}}_{G}:=\{x\in{\mathbb{R}}^{3}_{\geq 0}\,|\,x_{1}\leq 1/\alpha,\;x_{2}\leq 1/(\alpha\beta),\;x_{3}\leq 1/(\alpha\beta\gamma)\}.$
The system (19) admits a unique equilibrium $e=\begin{bmatrix}e_{1}&e_{2}&e_{3}\end{bmatrix}^{\top}$, where $e_{3}$ is the unique real and positive root of the polynomial
$Q(s):=\alpha\beta\gamma s^{m+1}+\alpha\beta\gamma s-1,$ (20)
$e_{1}=\beta\gamma e_{3}$, and $e_{2}=\gamma e_{3}$. Since ${{\mathcal{B}}_{G}}$ is a compact, convex, and invariant set, $e\in{\mathcal{B}}_{G}$.
If $e$ is locally asymptotically stable, then we may expect that all solutions converge to $e$. Tyson [33] proved that the system (19) admits a periodic solution whenever $e$ is unstable, but provided no information on the stability of the periodic solution. Indeed, he states [33, p. 312]: “Notice that we are not proving that this closed orbit is a global attractor of the torus. Though we might expect this from the computer simulations, it would be much more difficult to prove than simple existence”.
We now considerably strengthen the result of [33] by showing that, whenever $e$ is unstable, the system admits at least one periodic orbit that is attracting. We also have consistency of oscillations with respect to the initial conditions: the solution emanating from any $x(0)\in{{\mathcal{B}}_{G}}\setminus\{e\}$ converges to a periodic orbit.
The Jacobian of (19) is
$J(x)=\begin{bmatrix}-\alpha&0&-\frac{mx_{3}^{m-1}}{(1+x_{3}^{m})^{2}}\\ 1&-\beta&0\\ 0&1&-\gamma\end{bmatrix},$
so the system is 2-cooperative on ${\mathcal{B}}_{G}$, as $J(x)$ has the sign pattern (2) . We now show that it is also strongly $2$-cooperative. Using (19), it is easy to verify that if $x(t)$ is a solution of (19) with $x(0)\in{\mathcal{B}}_{G}$ and $x_{3}(0)=0$, then there exists some $\delta>0$ such that $t\in(t_{0},t_{0}+\delta)\implies x_{3}(t)>0$. In particular, the set $\left\{t\geq 0\ |\ x_{3}(t)=0\right\}$ is at most countable, and this implies that $M(t)$ in (4), which is obtained from $J(x)$, is irreducible for almost all $t$.
The characteristic polynomial of $J(e)$ is
$\det(sI_{3}-J(e))$ $=s^{3}+(\alpha+\beta+\gamma)s^{2}+(\alpha\beta+\alpha\gamma+\beta\gamma)s$
$+\alpha\beta\gamma+(me_{3}^{m-1}/(1+e_{3}^{m})^{2}),$
so the free coefficient is positive. Thus, Thm. 1 applies to this case study and yields the following corollary.
Corollary 1.
Consider the 3D Goodwin model (19) with equilibrium $e\in{\mathcal{B}}_{G}$. Suppose that $J(e)$ is unstable. Then, for any initial condition $x(0)\in{\mathcal{B}}_{G}\setminus\{e\}$, the solution of (19) converges to a periodic orbit.
The next numerical example demonstrates the result.
Example 2.
For the system (19) with $\alpha=0.5$, $\beta=0.4$, $\gamma=0.6$, and $m=10$, the box ${\mathcal{B}}_{G}$ becomes
${\mathcal{B}}_{G}=\{x\in{\mathbb{R}}^{3}_{\geq 0}\,|\,x_{1}\leq 2,\;x_{2}\leq 5,\;x_{3}\leq 25/3\},$
and the polynomial in (20) becomes
$Q(s)=0.12s^{11}+0.12s-1.$
The unique real and positive root of this polynomial is $e_{3}=1.1956$, so
$e=\begin{bmatrix}\beta\gamma e_{3}&\gamma e_{3}&e_{3}\end{bmatrix}^{\top}=\begin{bmatrix}0.2870&0.7174&1.1956\end{bmatrix}^{\top}.$
The characteristic polynomial is
$\det(sI_{3}-J(e))=s^{3}+1.5s^{2}+0.74s+1.1478,$
and applying the Routh stability criterion implies that $e$ is unstable. Indeed, the eigenvalues of $J(e)$ are $0.0062+\jmath 0.8711$, $0.0062-\jmath 0.8711$ and $-1.5125$, and correspond to the eigenvectors
$\begin{bmatrix}-0.2423-\jmath 0.5195\\ -0.5964+\jmath 0.0000\\ -0.3210+\jmath 0.4612\end{bmatrix},\;\begin{bmatrix}-0.2423+\jmath 0.5195\\ -0.5964-\jmath 0.0000\\ -0.3210-\jmath 0.4612\end{bmatrix},\;\begin{bmatrix}0.5999\\ -0.5393\\ 0.5910\end{bmatrix}.$
Fig. 2 depicts the solution $x(t)$ of (19) emanating from the initial condition $x(0)=\begin{bmatrix}0.1&0.1&0.1\end{bmatrix}^{\top}$. It can be seen that $x(t)$ converges to a periodic orbit.
Figure 2: Solution of the 3D Goodwin system in Example 2, with initial condition $x(0)=\begin{bmatrix}0.1&0.1&0.1\end{bmatrix}^{\top}$, converging to a periodic orbit. The equilibrium point $e$ is denoted by *.
IV-B Field-Noyes model
The Field-Noyes ODE model for the Belousov-Zhabotinskii reaction, derived in [11], is given by
$\begin{cases}\dot{x}_{1}&=s(x_{2}-x_{1}x_{2}+x_{1}-qx_{1}^{2}),\\ \dot{x}_{2}&=\frac{1}{s}(x_{3}f-x_{2}-x_{1}x_{2}),\\ \dot{x}_{3}&=w(x_{1}-x_{3}),\end{cases}$ (21)
where $s,q,f,w$ are positive constants. The state variables represent concentrations in chemical reactions, so the state space is $\Omega:={\mathbb{R}}^{3}_{\geq 0}$.
We assume throughout that $q\ll 1$ (the numerical parameter value in [11] is $q=8.375\times 10^{-6}$, see also [16]). Then, the closed box
[…]
with $y_{1}:=(1+q)^{-1}qf$ and $y_{2}:=(2q)^{-1}f$, is an invariant set of the dynamics (see [22]).
The system admits two equilibrium points in $\Omega={\mathbb{R}}^{3}_{\geq 0}$. The first is the origin (which is not in ${\mathcal{B}}_{FN}$). The second is $e=\begin{bmatrix}e_{1}&e_{2}&e_{3}\end{bmatrix}^{\top}\in{\mathcal{B}}_{FN}$, with
$e_{1}$ $=\frac{1}{2q}\left(1-f-q+\left((1-f-q)^{2}+4q(1+f)\right)^{1/2}\right),$
$e_{2}$ $=\frac{e_{1}f}{1+e_{1}}=(1+f-qe_{1})/2,$ (22)
$e_{3}$ $=e_{1}.$
The Jacobian of (21) is
$J(x)=\begin{bmatrix}s(1-x_{2}-2qx_{1})&s(1-x_{1})&0\\ -\frac{1}{s}x_{2}&-\frac{1}{s}(1+x_{1})&\frac{1}{s}f\\ w&0&-w\end{bmatrix},$
and has the sign pattern
$J(x)=\begin{bmatrix}*&<0&0\\ <0&*&>0\\ >0&0&*\end{bmatrix},\text{ for all }x\in\operatorname{int}({{\mathcal{B}}}),$
implying that (21) is strongly 2-cooperative, up to a coordinate transformation [36]. A calculation shows that the free coefficient of the polynomial $\det(\lambda I_{3}-J(e))$ is
$a_{0}$ $=w(2qe_{1}^{2}+(2q+f-1)e_{1}+2e_{2}-f-1),$
and using (IV-B) gives
$a_{0}$ $=we_{1}(2qe_{1}+q+f-1)$
$=we_{1}\left((1-f-q)^{2}+4q(1+f)\right)^{1/2}>0.$
Thus, Thm. 1 applies to this case study and yields the following corollary.
Corollary 2.
Consider the 3D Field-Noyes model (21) with equilibrium $e\in{\mathcal{B}}_{FN}$. Suppose that $J(e)$ is unstable. Then, for any initial condition $x(0)\in{\mathcal{B}}_{FN}\setminus\{e\}$, the solution of (21) converges to a periodic orbit.
Example 3.
For the system (21) with $q=9.374\cdot 10^{-6}$, $f=1$, $s=0.3$, and $w=0.2934$, the box ${\mathcal{B}}_{FN}$ becomes
[…]
where
$y_{1}=8.374\cdot 10^{-6},\ y_{2}=5.97\cdot 10^{4}.$
The corresponding equilibrium point is given by
$e=\begin{bmatrix}488.1780&0.9979&488.1780\end{bmatrix}^{\top}.$
The characteristic polynomial is
$\det(sI_{3}-J(e))=s^{3}+1630.8886s^{2}-4.8311s+1.1722,$
so the free coefficient is positive, and since one of the coefficients is negative, $J(e)$ is unstable. Fig. 3 depicts the solution $x(t)$ of (21) emanating from the initial condition $x(0)=\begin{bmatrix}732.2670&9.9795&732.2670\end{bmatrix}^{\top}$. It can be seen that $x(t)$ converges to a periodic orbit.
Figure 3: A solution of the Field-Noyes system in Example 3 converges to a periodic orbit.
V Discussion
Strongly 2-cooperative systems enjoy an important asymptotic property: all bounded solutions that keep a positive distance from the set of equilibria converge to a periodic orbit. It is important to find cases where an attracting periodic orbit can be shown to exist. Here, we derived a simple sufficient condition that guarantees the existence of an attracting periodic orbit for 3D systems. The proof relies on the asymptotic and spectral properties of strongly 2-cooperative systems. Our result provides a unifying treatment of interesting case studies in systems biology and allows to derive stronger theoretical results than those available in the literature, while simplifying and unifying the proofs.
An $n$-dimensional system that is $(n-1)$-cooperative is , up to a coordinate transformation, a competitive system [35]. In particular, $2$-cooperative 3D systems are competitive systems. Indeed, let $D:=\mathop{\rm diag}\nolimits(1,-1,1)$ and $P:=\begin{bmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{bmatrix}$. Then for $n=3$ we have $\bar{A}_{2}=\begin{bmatrix}*&\geq 0&\leq 0\\ \geq 0&*&\geq 0\\ \leq 0&\geq 0&*\end{bmatrix}$, and
$DP\bar{A}_{2}(DP)^{-1}=\begin{bmatrix}*&\leq 0&\leq 0\\ \leq 0&*&\leq 0\\ \leq 0&\leq 0&*\end{bmatrix}.$
In other words, if $\dot{x}=Ax$ is a $2$-cooperative 3D system and $y(t):=DPx(t)$ then the $y$-system $\dot{y}=DPA(DP)^{-1}y$ is competitive. Thus, our results may also be interpreted as providing a sufficient condition guaranteeing that 3D competitive systems admit an attracting periodic orbit.
An interesting topic for further research is to find conditions that guarantee uniqueness of the periodic orbit. Then, our result would imply that the periodic orbit is a global attractor for solutions emanating from the invariant set ${\mathcal{H}}$. Another interesting research direction is to extend our result to systems with dimension $n>3$. Finally, our results may find application to the design of oscillators, which attracts considerable interest e.g. in synthetic biology [6, 23, 24].